# Simultaneous Source Separation, Synchronization, Localization and Mapping for 6G Systems

### v1

@@ V-3 Selected SPA Messages
The messages $\xi\big(\overline{a}^{(i)}_{m,n}\big)$ sent from the factor node $\overline{q}_{\text{BS}}\big(\overline{\bm{p}}^{(i)}_{\text{va},m},\overline{r}^{(i)}_{m,n}$, $\overline{a}^{(i)}_{m,n},\bm{x}_{n}^{(i)};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ to the variable nodes corresponding to the measurement-oriented association variables $\overline{a}^{(i)}_{m,n}$ are given by
@@
$\hskip 56.9055pt\overline{a}^{(i)}_{m,n},\bm{x}_{n}^{(i)},\overline{j}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\,\mathrm{d}\bm{x}_{n}^{(i)}\hskip 0.85358pt\mathrm{d}\overline{\bm{p}}^{(i)}_{\text{va},m}$ (21)
where the new PVA state $\overline{\bm{\mathsfbr{y}}}^{(i)}_{m,n}=[\overline{\bm{p}}^{(i)}_{\text{va},m}\hskip 2.84526pt\overline{\bm{r}}^{(i)}_{m,n}]^{\text{T}}$ is replaced by its entries. Using the expression of $\overline{q}_{\text{BS}}\big(\overline{\bm{p}}^{(i)}_{\text{va},m}$, $\overline{r}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n},\bm{x}_{n}^{(i)};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ in (19), Eq. (21) simplifies to $\xi\big(\overline{a}^{(i)}_{m,n}\big)\!=\!1$ for $\overline{a}^{(i)}_{m,n}\!\in\!{\cal{K}}_{n}$, and for $\overline{a}^{(i)}_{m,n}\!=\hskip-0.85358pt0$ it becomes
@@
$\hskip 28.45274pt\times f({\bm{z}_{\text{bs}}}^{(i)}_{n}|\bm{x}_{n}^{(i)},\overline{\bm{p}}^{(j,i)}_{\text{va},m})\,\tilde{f}^{(i)}(\bm{x}_{n}^{(i)})\hskip 0.85358pt\mathrm{d}\bm{x}_{n}^{(i)}\hskip 0.85358pt\mathrm{d}\overline{\bm{p}}^{(j,i)}_{\text{va},m}\hskip 0.85358pt$. (22)
@@ VI Results
We consider a simple indoor scenario as shown in Fig. 2. The scenario comprises four BSs and 4 reflective surfaces, i.e., $4$ VAs per BS. Three MTs move along tracks which are observed for $400$ time instances $n$ with observation period $\Delta T=1\,$s. Measurements are generated according to the system model in Section IV. The signal-to-noise-ratio (SNR) is set to $40\,\mathrm{dB}$ at a LOS distance of 11\,m. The amplitudes of the MPCs (including the LOS component) are calculated using free-space path loss and an additional attenuation of $3\,\mathrm{dB}$ for each reflection at a surface. The detection threshold is $\gamma=8\,$dB and $p_{\text{d}}=0.8$. False positive measurements are generated according to the model in Section III with a mean number of FPs $\mu_{\text{fp}}=5$. For the calculation of the measurement variances, we assume a system bandwidth of $B=100\,\mathrm{MHz}$ at a center frequency of $f_{\mathrm{c}}=6\,\mathrm{GHz}$. The arrays employed at MTs and PAs are of identical geometry with $H=H^{(j)}=4$ antenna elements forming a uniform rectangular array spaced at $\lambda/2$, where $\lambda=c/f_{\mathrm{c}}$ is the carrier wavelength. The true BS biases are equally spread in a range of $\pm 30$ m ($\pm 100$ ns), while the MT biases are spread in a small range between $\pm 0.5$ m and the true delay biases drift over time following a zig-zag pattern shown in Figure 5 with a slope of $0.01$ ns. We use $10^{4}$ particles. The particles for the initial MT position and velocity are drawn from i.i.d. Gaussian distributions with center being drawn from the same distribution for each realization, which is centered around the true MT state, and initialization standard deviations given as $0.5~\text{m}$ and $0.1~\text{m/s}$. The initial orientation is determined from the velocity, assuming the agent to move in forward-facing direction. When $n\hskip-0.85358pt=\hskip-0.85358pt0$, the number of VAs is 0, i.e., no prior map information is available. The prior for new PVA states $f_{\text{n}}\big(\overline{\bm{x}}^{(j)}_{m,n}|\bm{x}_{n}\big)$ is uniform on the square region given by [45 m,45 m]×[45 m,45 m][-$\text$$\text{45 m}$$\times$$\text$$\text{45 m}$] around the center of the floor plan shown in Fig. 2 and the mean number of new PVAs at time $n$ is $\mu_{\text{n}}=0.01$. As a proposal distribution for the prior we use anulus shaped distributions, centered around the measurement for numerical efficiency. The probability of survival is $p_{\mathrm{s}}=0.999$. The confirmation threshold and the pruning threshold, are given as $p_{\mathrm{cf}}=0.5$ and $p_{\mathrm{pr}}=10^{-3}$, respectively. The BS synchronization biases initialized uniformly distributed between $100~\mathrm{ns}$ and $100~\mathrm{ns}$ and the MT synchronization biases are initialized uniformly distributed between $1~\mathrm{ns}$ and $1~\mathrm{ns}$. The MT state transition PDF $f(\bm{x}_{n}|\bm{x}_{n-1})$ is modeled independently for MT bias, MT orientation and the MTs position-velocity state $\bm{\mathsfbr{p}}_{n},\bm{\mathsfbr{v}}_{n}]$. The MT position-velocity state is modeled by a constant-velocity and stochastic-acceleration model [29] with acceleration variances set to $\sigma_{\text{w}}=10^{-3}\,\mathrm{m/s^{2}}$. The orientation state transition is modeled by a Gaussian random walk with standard deviation of $3^{\circ}$. The state transitions of both, the MT biases and the BS biases, are modeled by a Gaussian random walk with driving noise set to $\sigma_{b\,BS}$ = $\sigma_{b\,MT}$ = $0.1~\text{ns}$.
@@
The performance is measured in terms of the root mean squared error (RMSE) of the MT position as well as the optimal subpattern assignment (OSPA) error [30] of all VAs with cutoff parameter set to $1~\mathrm{m}$ and order set to $2$. The mean optimal subpattern assignment (MOSPA) errors and RMSEs are obtained using $100$ simulation runs.
@@
Figure 4: Factor graph corresponding to the factorization shown in (15). Dashed arrows represent messages that are only passed in one direction. The detailed graph in (a) represents the green subgraphs of the overall factor graph in (b). The following short notation is used. Variables and variable nodes: $\bm{x}_{i}\triangleq\bm{x}^{(i)}_{n}$, $\underline{\bm{y}}^{i}_{k}\triangleq\underline{\bm{y}}^{(i)}_{k,n}$, $\overline{\bm{y}}^{i}_{m}\triangleq\overline{\bm{y}}^{(i)}_{m,n}$, $\underline{a}_{k}\triangleq\underline{a}^{(i)}_{k}$, $\overline{a}_{m}\triangleq\overline{a}^{(i)}_{m}$, $b_{\text{bs}}^{j}\triangleq b_{\text{bs},n}^{(j)}$, $\overline{j}^{(i)}_{m,n}\triangleq j_{m}$, $K^{-}\triangleq K_{n-1}$, and $K\triangleq K_{n}$ with $K_{n}=K_{n}^{(I)}+M^{(I)}$. Factor node: $f_{i}\triangleq f(\bm{x}^{(i)}_{n}|\bm{x}^{(i)}_{n-1})$, $f^{i}_{k}\triangleq f(\bm{y}^{(j)}_{k}|\bm{y}^{(j)-}_{k})$, $f^{j}_{\text{b}}\triangleq f(b_{\text{bs},n}^{(j)}|b_{\text{bs},n-1}^{(j)})$, and $\psi_{k,m}\triangleq\Psi(\underline{a}^{(i)}_{k,n},\overline{a}^{(i)}_{m,n})$. Messages: $q_{\text{x}}^{i}\triangleq f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $q_{\text{b}}^{j}\triangleq f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, and $q_{k}\triangleq f(\bm{y}_{k}|\bm{z}_{1:n})$, where $\bm{y}_{k}$ represents an entry of $\bm{\mathsfbr{y}}_{n}\!=\hskip-0.85358pt\big[\underline{\bm{\mathsfbr{y}}}^{(I)\hskip 0.85358pt\text{T}}_{n}\hskip 0.85358pt\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(I)\text{T}}_{n}\big]^{\text{T}}$. The dashed arrows indicate messages representing MT and PVA beliefs of time n1n-1 and $n+1$, which are only propagated forward in time and the minus sign “-” indicates beliefs from the previous time n1n-1.
@@
To investigate the proposed method, we evaluate four algorithm variants: (i) unsynchronized and interfering BSs, (ii) synchronized and interfering BSs, (iii) unsynchronized and orthogonal BSs, and (iv) synchronized and orthogonal BSs, with the (iv) being the reference case that corresponds to state-of-the-art (SOTA) cooperative MP-SLAM according to [17]. Here, “unsynchronized” refers to unknown delay biases between BSs, while “interfering” indicates that BSs use identical transmission sequences and, thus, it is unknown which MPC and, thus, which PVA in the received radio signal corrresponds to which BS. For unsychronized and interfering BS, Figure 2 shows the true PVAs and MT positions along with their respective MMSE estimates of the proposed algorithm for a single realization. One can obtain that the PVA estimates converge to the true colors, which is possible – although there is no prior information about the delay biases – since the AOD measurements depend on the (known) BS orientation (see Section III). Note that without AOD measurements the proposed algorithm could still differentiate, which VA belongs to which BS, but the colors would be different to the ground truth. The mean RMSE of BS biases and MT biases for $n\in\{101,\cdots,400\}$ was $26~\mathrm{ps}$ and $24~\mathrm{ps}$, respectively. The mean error of MT and BS biases are illustrated in Figure 5. Figure 6 provides the numerical results for the experiment in terms of (a) the cumulative frequency and (b) average RMSE of the MT position error, (c) PVA MOSPA, and (d) mean number of PA association errors per time $n$. One can observe that the proposed algorithm takes several times $n$ to fully converge, leading to an increased MT position error and PVA MOSPA for the first $100$ steps $n$, but then reaches the performance of the reference algorithm (synchronized, orthogonal BSs). For the investigated setup, there are no PA association errors at any time $n$ and thus, there is no significant performance loss for unsynchronized BSs.
@@ VII Conclusion
This paper presented a Bayesian factor-graph framework for cooperative MP-SLAM with unsynchronized and interfering base stations. The proposed method jointly performs synchronization, source separation, and environment mapping via sequential message passing on a structured SPA. Simulation results show no significant performance degradation of the proposed algorithm compared to state-of-the-art cooperative MP-SLAM that assumes synchronized BSs and no inter-BS interference. The SPA framework is flexible and can incorporate additional sensing modalities such as IMU or radar, enabling seamless integration into future 6G ISAC systems. Future work will address real-world experiments with 5G PRS data, extensions to dynamic network scenarios or to integrate particle flow to improve the birth of new features through the highly nonlinear system models [31, 32].

### v2

@@ V-3 Selected SPA Messages
The messages $\xi\big(\overline{a}^{(i)}_{m,n}\big)$ sent from the factor node $\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},\bm{b}^{i}_{\text{bs},n},$ $\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ to the variable nodes corresponding to the measurement-oriented association variables $\overline{a}^{(i)}_{m,n}$ are given by
@@
$\hskip 42.67912pt\bm{x}_{n}^{(i)},\bm{b}^{i}_{\text{bs},n},\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\,\mathrm{d}\bm{x}_{n}^{(i)}\hskip 0.85358pt\mathrm{d}\overline{\bm{p}}^{(i)}_{\text{va},m}$ (22)
where the new PVA state $\overline{\bm{\mathsfbr{y}}}^{(i)}_{m,n}=[\overline{\bm{p}}^{(i)}_{\text{va},m}\hskip 2.84526pt\overline{\bm{r}}^{(i)}_{m,n}]^{\text{T}}$ is replaced by its entries. Using the expression of $\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},\bm{b}^{i}_{\text{bs},n},$ $\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ in (20), Eq. (22) simplifies to $\xi\big(\overline{a}^{(i)}_{m,n}\big)\!=\!1$ for $\overline{a}^{(i)}_{m,n}\!\in\!{\cal{K}}_{n}$, and for $\overline{a}^{(i)}_{m,n}\!=\hskip-0.85358pt0$ it becomes
@@
$\hskip 0.0pt\times f({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\hskip 0.85358pt|\bm{x}_{n}^{(i)},\overline{\bm{p}}^{(j,i)}_{\text{va},m},{b}^{(j,i)}_{\text{bs},n})\,\tilde{f}^{(i)}(\bm{x}_{n}^{(i)})\hskip 0.85358pt\mathrm{d}\bm{x}_{n}^{(i)}\hskip 0.85358pt\mathrm{d}\overline{\bm{p}}^{(j,i)}_{\text{va},m}\hskip 0.85358pt$. (23)
@@ VI Results
We consider a simple indoor scenario as shown in Fig. 2. The scenario comprises four BSs and 4 reflective surfaces, i.e., $4$ VAs per BS. Three MTs move along tracks which are observed for $400$ time instances $n$ with observation period $\Delta T=1\,$s. Measurements are generated according to the system model in Section IV. The signal-to-noise-ratio (SNR) is set to $40\,\mathrm{dB}$ at a LOS distance of 11\,m. The amplitudes of the MPCs (including the LOS component) are calculated using free-space path loss and an additional attenuation of $3\,\mathrm{dB}$ for each reflection at a surface. The detection threshold is $\gamma=8\,$dB and $p_{\text{d}}=0.8$. False positive measurements are generated according to the model in Section III with a mean number of FPs $\mu_{\text{fp}}=5$. For the calculation of the measurement variances, we assume a system bandwidth of $B=100\,\mathrm{MHz}$ at a center frequency of $f_{\mathrm{c}}=6\,\mathrm{GHz}$. The arrays employed at MTs and PAs are of identical geometry with $H=H^{(j)}=4$ antenna elements forming a uniform rectangular array spaced at $\lambda/2$, where $\lambda=c/f_{\mathrm{c}}$ is the carrier wavelength. The true BS biases are equally spread in a range of $\pm 30$ m ($\pm 100$ ns), while the MT biases are spread in a small range between $\pm 0.5$ m and the true delay biases drift over time following a zig-zag pattern with a slope of $0.01$ ns. We use $10^{4}$ particles. The particles for the initial MT position and velocity are drawn from i.i.d. Gaussian distributions with center being drawn from the same distribution for each realization, which is centered around the true MT state, and initialization standard deviations given as $0.5~\text{m}$ and $0.1~\text{m/s}$. The initial orientation is determined from the velocity, assuming the agent to move in forward-facing direction. When $n\hskip-0.85358pt=\hskip-0.85358pt0$, the number of VAs is 0, i.e., no prior map information is available. The prior for new PVA states $f_{\text{n}}\big(\overline{\bm{x}}^{(j)}_{m,n}|\bm{x}_{n}\big)$ is uniform on the square region given by [45 m,45 m]×[45 m,45 m][-$\text$$\text{45 m}$$\times$$\text$$\text{45 m}$] around the center of the floor plan shown in Fig. 2 and the mean number of new PVAs at time $n$ is $\mu_{\text{n}}=0.01$. As a proposal distribution for the prior we use anulus shaped distributions, centered around the measurement for numerical efficiency. The probability of survival is $p_{\mathrm{s}}=0.999$. The confirmation threshold and the pruning threshold, are given as $p_{\mathrm{cf}}=0.5$ and $p_{\mathrm{pr}}=10^{-3}$, respectively. The BS synchronization biases initialized uniformly distributed between $100~\mathrm{ns}$ and $100~\mathrm{ns}$ and the MT synchronization biases are initialized uniformly distributed between $1~\mathrm{ns}$ and $1~\mathrm{ns}$. The MT state transition PDF $f(\bm{x}_{n}|\bm{x}_{n-1})$ is modeled independently for MT bias, MT orientation and the MTs position-velocity state $\bm{\mathsfbr{p}}_{n},\bm{\mathsfbr{v}}_{n}]$. The MT position-velocity state is modeled by a constant-velocity and stochastic-acceleration model [29] with acceleration variances set to $\sigma_{\text{w}}=10^{-3}\,\mathrm{m/s^{2}}$. The orientation state transition is modeled by a Gaussian random walk with standard deviation of $3^{\circ}$. The state transitions of both, the MT biases and the BS biases, are modeled by a Gaussian random walk with driving noise set to $\sigma_{b\,BS}$ = $\sigma_{b\,MT}$ = $0.1~\text{ns}$.
@@
The localization and mapping performance is measured in terms of the mean root mean squared error (RMSE) of the MT positions as well as the optimal subpattern assignment (OSPA) error [30] of all VAs with cutoff parameter set to $1~\mathrm{m}$ and order set to $2$. The mean optimal subpattern assignment (MOSPA) errors and RMSEs are obtained using $100$ simulation runs.
@@
Figure 4: Factor graph corresponding to the factorization shown in (16). Dashed arrows represent messages that are only passed in one direction. The detailed graph in (a) represents the green subgraphs of the overall factor graph in (b). The following short notation is used. Variables and variable nodes: $\bm{x}_{i}\triangleq\bm{x}^{(i)}_{n}$, $\underline{\bm{y}}^{i}_{k}\triangleq\underline{\bm{y}}^{(i)}_{k,n}$, $\overline{\bm{y}}^{j}_{m}\triangleq\overline{\bm{y}}^{(j,i)}_{m,n}$, $\underline{a}_{k}\triangleq\underline{a}^{(i)}_{k}$, $\overline{a}_{m}\triangleq\overline{a}^{(i)}_{m}$, $b_{\text{bs}}^{j}\triangleq b_{\text{bs},n}^{(j)}$, $\overline{j}_{m}\triangleq\overline{j}^{(i)}_{m,n}$, $^{-}\triangleq K_{n-1}, KK$$\triangleq K^{(i)}_{n}$ $\tilde{K}=J+K^{(i)}_{n}$, and $K^{+}\triangleq K_{n}$. Factor node: $f_{i}\triangleq f(\bm{x}^{(i)}_{n}|\bm{x}^{(i)}_{n-1})$, $f_{y,k}\triangleq f(\underline{\bm{y}}_{k,n}|\bm{y}_{k,n-1})$ for $i=1$, $f_{y,k}\triangleq f(\underline{\bm{y}}^{(i)}_{k,n}|\bm{y}^{(i-1)}_{k,n-1})$ for $i>1$, $f^{j}_{\text{b}}\triangleq f^{(i)}(b_{\text{bs},n}^{(j,i)}|b_{\text{bs},n}^{(j,i-1)})$, and $\psi_{k,m}\triangleq\Psi(\underline{a}^{(i)}_{k,n},\overline{a}^{(i)}_{m,n})$. Messages: $\tilde{f}_{\text{x}}^{i}\triangleq f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $\tilde{f}_{\text{b}}^{j}\triangleq f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, and $\tilde{f}_{k}\triangleq f(\bm{y}_{k}|\bm{z}_{1:n})$, where $\bm{y}_{k}$ represents an entry of $\bm{\mathsfbr{y}}_{n}\!=\hskip-0.85358pt\big[\underline{\bm{\mathsfbr{y}}}^{(I)\hskip 0.85358pt\text{T}}_{n}\hskip 0.85358pt\hskip 0.85358pt\overline{\bm{\mathsfbr{y}}}^{(I)\text{T}}_{n}\big]^{\text{T}}$. The dashed arrows indicate messages representing MT and PVA beliefs of time n1n-1 and $n+1$, which are only propagated forward in time and the minus sign “-” indicates beliefs from the previous time n1n-1.
@@
To investigate the proposed method, we evaluate four algorithm variants: (i) unsynchronized and interfering BSs, (ii) synchronized and interfering BSs, (iii) unsynchronized and orthogonal BSs, and (iv) synchronized and orthogonal BSs, with the (iv) being the reference case that corresponds to state-of-the-art (SOTA) cooperative MP-SLAM according to [17]. Here, “unsynchronized” refers to unknown delay biases between BSs, while “interfering” indicates that the BSs use identical transmission sequences and, thus, it is unknown which MPC in the received radio signal (and, hence, which PVA) corrresponds to which BS. For unsychronized and interfering BS, Figure 2 shows the true PVAs and MT positions along with their respective MMSE estimates of the proposed algorithm for a single realization. One can obtain that the PVA estimates converge to the true colors, which is possible although there is no prior information about the delay biases since the AOD measurements depend on the (known) BS orientation (see Section III). Note that without AOD measurements the proposed algorithm could still differentiate, which VA belongs to which BS, but the colors would be different to the ground truth. The mean RMSE of BS biases and MT biases for $n\in\{101,\cdots,400\}$ was $26~\mathrm{ps}$ and $24~\mathrm{ps}$, respectively. The mean error of MT and BS biases are illustrated in Figure 5. Figure 6 provides the numerical results for the experiment in terms of (a) the cumulative frequency and (b) average RMSE of the MT position error, (c) PVA MOSPA, and (d) mean number of PA association errors per time $n$. One can observe that the proposed algorithm takes several times $n$ to fully converge, leading to an increased MT position error and PVA MOSPA for the first $100$ steps $n$, but then reaches the performance of the reference algorithm (synchronized, orthogonal BSs). For the investigated setup, there are no PA association errors at any time $n$ and thus, there is no significant performance loss for unsynchronized BSs.
@@ VII Conclusion
This paper presented a Bayesian factor-graph framework for cooperative MP-SLAM with unsynchronized and interfering base stations. The proposed method jointly performs synchronization, source separation, and environment mapping via sequential message passing on a structured SPA. Simulation results show no significant performance degradation of the proposed algorithm compared to state-of-the-art cooperative MP-SLAM that assumes synchronized BSs and no inter-BS interference. The proposed BS-dependent data association model constitutes a principled approach for classifying features by arbitrary properties, such as reflection order or feature type (scatterers versus walls). The SPA framework is flexible and can incorporate additional sensing modalities such as IMU or radar, enabling seamless integration into future 6G ISAC systems. Future work will address real-world experiments with 5G PRS data, extensions to dynamic network scenarios or to integrate particle flow to improve the birth of new features through the highly nonlinear system models [31, 32].